OpenStudy (anonymous):

integrate(xcoshx, x, 0, 1)

OpenStudy (sirm3d):

\[\int_0^1x\cosh x \;\mathrm dx\] try integration by parts

OpenStudy (anonymous):

I did. But the answer for the integral= (e-1)/e. And I can get that.

OpenStudy (anonymous):

*cant

OpenStudy (anonymous):

I got to [(1)sinh(1)-(0)sinh(0)]-[cosht]^1(a)0(b)

OpenStudy (anonymous):

the 1 and 0 are the a and b for the integration of cosh t

OpenStudy (sirm3d):

\[\int x\cosh x \mathrm dx=x\sinh x - \cosh x\] \[\sinh x = (1/2) (e^x - e^{-x})\\\cosh x = (1/2)(e^x+e^{-x})\] so sinh 1 = (1/2)e^1-e^(-1)